Two cans x and y both contain some water. From x Jim pours as much water into y as y already contains. Then from y he pours as much water into x as y presently has. Finally he pours from x to y as much water as y presently has. Each can now contains 24 ounces of water. how many ounces of water were in each can at the start?
Answer by psbhowmick(878)

Let the cans initially contain 'u' and 'v' amounts of water respectively.

As Jim pours ,from X, as much water into Y as Y already contains so after the first operation, X contains (u-v) and Y contains (v + v =) '2v' amount of water.

Next, he pours 'v' amount of water from Y into X (as presently Y has 'v' amount of water).
So volume of water in Y becomes 'v' and amount of water in X becomes (u-v) + v = u.

Again, Jim pours 'v' amount of water into Y (as presently Y has 'v' amount of water).
So volume of water in Y becomes '2v' and amount of water in X becomes (u-v).
These are the final amounts of water present in X and Y.
Final content in each vessel is 24 ounces.

Thus we have two equations:
u - v = 24 and 2v = 24.
The equation: 2v =24 gives v = 24/2 = 12.
Then from u - v = 24 we get u = 24 + v = 24 + 12 = 36.

As the initial contents of the vessels were 'u' and 'v' so they contained 36 and 12 ounces of water respectively.